OpenStudy (jaredstone4):

Solve the homogenous differential equation for y: y' = (5x+9y) /x I got it down to xdv / (5+8v) = dx, but I don't know where to go from here.

OpenStudy (anonymous):

move x to the other side it becomes separable

OpenStudy (jaredstone4):

so it'd be dv/(5+8v) = dx/x ?

OpenStudy (anonymous):

yea

OpenStudy (jaredstone4):

Okay so I did that, integrated, and now I have ln l5+8vl = ln(Cx). From here?